A radio station broadcast at a frequency of 600KHZ. If the speed of light in air is 3 × 108ms−1−1. Calculate the wavelength of the radio wave
1
Expert's answer
2019-09-10T13:48:14-0400

The wavelength of the radio wave


"\lambda=\frac{c}{\nu}=\frac{3\times 10^8\:\rm m/s}{600\times 10^3\:\rm Hz}=500\:\rm m"